OpenStudy (anonymous):

The radius of a cylinder is 15 cm, and its volume, found using 3.14 to approximate pi, is 19,782 cm^3. What is the height of the cylinder? ____cm

OpenStudy (anonymous):

Formula: \(V = \pi r^2 h\) Plug in what we know: \(19782 = (3.14)(15^2)h\) \(19782 = (3.14)(225)h\) What's 225 * 3.14?

OpenStudy (anonymous):

Is that the answer?

OpenStudy (anonymous):

No, just multiply 225 * 3.14

OpenStudy (anonymous):

\(225 \times 3.14\)

OpenStudy (anonymous):

800.7

OpenStudy (anonymous):

No, check again

OpenStudy (anonymous):

706.5

OpenStudy (anonymous):

Yes, now we divide: 19782 / 706.5

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

28?

OpenStudy (anonymous):

Yes, that is our answer. So the height is 28.

OpenStudy (anonymous):

Thank You so much